# Topological Devil’s staircase in a constrained kagome Ising antiferromagnet

### v1

@@ Abstract
Two famous examples of unusual phase transitions are particularly relevant here. The first is the Kasteleyn transition, originally described by Kasteleyn in 1963 [7] in the context of dimer models [8, 9]. It occurs in constrained models [10, 11, 12] where the only defects that respect the constraints span the whole lattice; thus, their energy grows linearly with the size of the system. As a consequence, they are completely absent up to the temperature at which their free energy becomes negative due to their entropic contribution. The transition is completely asymmetric, with a specific heat that only diverges on the high-temperature side.
In frustrated models, the Kasteleyn transition may appear in more or less exotic forms, when constraints are imposed by infinite coupling constants. It is well known to play a key role in kagome ice [13, 14, 15], as well as in spin ice under a [100] field [16, 17, 17, 18, 19]. A case of direct relevance for the present study is the antiferromagnetic Ising model on the triangular lattice [20]. When first and second neighbor couplings are included, the ground state consists of alternating lines of up- and down spins. It breaks the $\mathbb{Z}_{3}$ rotational lattice symmetry, as well as the $\mathbb{Z}_{2}$ symmetry of exchanging spins up and down. For that model, both symmetries are restored through a first-order transition. However, if a fifth-neighbor coupling is included, the transition is a two-step process, and in the constrained limit of infinite first-neighbor coupling, the $\mathbb{Z}_{2}$ symmetry is first restored through a Kasteleyn transition before the $\mathbb{Z}_{3}$ symmetry is restored through a first-order transition.
@@
In this Letter, we study the antiferromagnetic Ising model on the kagome lattice with up to third neighbor couplings (Eq. 1) in a phase that breaks the rotational $\mathbb{Z}_{3}$ symmetry and the $\mathbb{Z}_{2}$ up-down symmetry. In the appropriate constrained limit of infinite first and third neighbor couplings, where the only allowed defects are system-spanning, we show that the transition is a highly exotic form of a Kasteleyn mechanism giving rise to a topological devil’s staircase  [^1] Here, we refer to the physical notion of a complete devil’s staircase [22], where the locked portions of the function fill the whole range of temperatures (up to the final critical point at $T=4J_{2}$). Although linear defects are completely suppressed at low temperature, as in the Kasteleyn transition, their density, instead of growing as $(T-T_{c})^{1/2}$, grows through an infinite and increasingly dense sequence of jumps. By contrast to the standard devil’s staircase, this behavior is not related to commensurate steps, but is controlled by an integer index of topological origin, the ratio of the densities of linear defects and of zero-energy domain walls.
Figure 1: Spin configurations of the kagome lattice Ising antiferromagnet Eq. (1) are represented by blue and red dots. Orange segments show an alternative representation. (a) Main family of strings ground states : domains of left- and right-pointing orange arrows are delimited by A and B strings. (b) Zero-energy double domain walls (DDWs) correspond to arrows rotated all in the same direction and are entropically suppressed in the thermodynamic limit. (c) C-string excitations in the first plateau are characterized by different orientations of the arrows inside the DDWs, which introduce an energy cost. The density of C-strings ($n_{C}$) is equal to the density of ferromagnetic nearest-neighbors on a horizontal dense row.
Model and partially ordered ground states. In this Letter, we will mostly focus on the properties of the constrained model that corresponds to $J_{1}$ and $J_{3}$ infinite in the Hamiltonian:
@@
where $\langle i,j\rangle_{n}$ denote pairs of $n^{\mathrm{th}}$ nearest-neighbors on the kagome lattice, and $\sigma_{i}=\pm 1$. In contrast to already extensively studied limits where the third-neighbor interaction is only across the hexagon [25, 26, 27, 28, 29, 30] or along the nearest-neighbor bonds [31, 32, 33, 34, 35, 36, 37]  [^2] A notable case is the SM of Ref. [64], which studies the $J_{2}=0$ limit with both 3rd neighbor couplings being non-zero., here, the third-neighbor interaction is set to the same value in both directions, forming triangular sublattices (Fig. 1a). For finite $J_{1}$ and $J_{3}$, the range $0<J_{2}\leq J_{3}<J_{1}$ corresponds to a single phase of ground state energy per site $E/N=-2J_{1}/3+2J_{2}/3-J_{3}/3$, referred to as the strings phase in Ref. [39]. In that phase, the macroscopic degeneracy of the nearest-neighbor model is only partially lifted, yielding a residual entropy evaluated numerically to 1/3 of that of the nearest-neighbor Ising antiferromagnet on the triangular lattice.
@@
Indeed, this entropy is compatible with a simple family of ground states, illustrated in Fig. 1(a). Seeing the kagome lattice as a stacking of dense and sparse rows, the dense rows form a rectangular lattice. Consider a configuration on this rectangular lattice where the order is antiferromagnetic in the direction of the dense rows and ferromagnetic in the perpendicular direction. The energy minimum is reached if the remaining spins, which live on a $J_{3}$ triangular sub-lattice, are in the ground state of that model, with one frustrated bond on each $J_{3}$ triangle. Since the number of these spins is 1/3 of the total number of spins, the residual entropy of this family is 1/3 of that of the triangular lattice. To visualize these states (Fig. 1(a)), it is convenient to draw dual-lattice bonds through the unsatisfied $J_{1}$ bonds. In the ground state, these dual bonds must not appear isolated; instead, they must belong to a pair of bonds meeting at 60º in the center of a hexagon, forming an arrow. If we define domains as contiguous regions where all arrows point in the same direction, the partially ordered family can be constructed by partitioning the lattice into domains with antiparallel directions  [^3] This is similar in spirit to the construction of the triangular lattice Ising antiferromagnetic ground states starting from the stripe order and introducing double domain walls, see e.g. [20, 65] and references therein.. The boundaries between the uniform regions form domain walls living on the bonds of the triangular lattice and connecting the centers of the hexagons. We call these domain walls A or B lines if they are surrounded by converging or diverging arrows, respectively (see EM1), or equivalently if they go through crossed or empty hexagons. In order to minimize $J_{3}$ interactions, the A and B lines may never propagate in parallel to the arrow directions or intersect, which forces them to form system-spanning zero energy domain walls. Note that the number of A and B lines may at most differ by one since every domain must be bounded by an A line on one side and a B line on the other. In the thermodynamic limit, their densities must therefore be equal.
@@
In the partially ordered family of states, the trajectories of A and B lines determine the spin configuration on the disordered $J_{3}$ triangular sub-lattice, where they separate antiferromagnetically aligned spins in the same row. Since adjacent spins on sparse rows have a $2/3$ probability of being antiferromagnetically aligned, the average density of A and B strings in the thermodynamic limit is $1/3$.
@@
Figure 2: Numerical results from CTMRG for $\chi=100$ suggest a topological devil’s staircase. Only temperatures where the results are converged in bond dimension are shown. (a) Energy. (b) Ratio of the strings density, revealing well-defined, integer-valued plateaus. (c) C-strings density $n_{C}$, related to the $\mathbb{Z}_{3}$ order parameter (see EM3). A careful inspection reveals a continuous temperature dependence in the apparent plateaus of $n_{C}$ [41]
@@
Interestingly, there is another family of zero energy extended defects, namely double domain walls (DDW) formed by replacing a B line with any number of arrows tilted by 60 or 120 degrees, as shown in Fig. 1(b). Such a DDW creates a defect in the antiferromagnetic order of the dense rows, and if a finite density of them (with random positions) was present in the ground state the $\mathbb{Z}_{2}$ symmetry would not be broken. These DDWs may meander in the same way that A and B lines do, but they have a thickness of one lattice unit. Thus, the entropy of configurations with one DDW is equivalent to the entropy of configurations with only A and B lines in a smaller system where one column has been removed. Since entropy is extensive, removing a column has an entropic cost linear in system size, and therefore DDWs are completely suppressed in the thermodynamic limit, inducing an entropy-driven partial order [42, 43]. This is in agreement with the residual entropy being exactly that of the simple family of states described above. There are also ground states which cannot be described by replacing B lines with DDW’s. Their number grows sub-extensively with system size and they also do not contribute to macroscopic properties of the ground state [41].
@@
Excitations in the constrained limit. Let us discuss more specifically the constrained limit of infinite $J_{1}$ and $J_{3}$. In that limit, the only possible excitations correspond to deviations from the $J_{2}$ ground state rule, which imposes a single bond to be unsatisfied in each $J_{2}$ triangle. Such excitations are only possible if a DDW is present. They are rooted in the development of domains with changes of the orientation of the arrows inside the DDW, possibly accompanied by portions of B-lines which decorate the DDW (see panel (c) of Fig. 1). From now on, these DDW with possible changes of orientation and decorations will be called C-lines. The energy cost of such C-lines comes from the presence of two types of local defects: (i) hexagons with two arrows that form a “K” instead of a cross, with energy $34J_{2}/$ 3, and (ii) hexagons with three arrows forming a “star”, with energy $4J_{2}$. As in the Kasteleyn transition, excitations with a finite density of defects are completely suppressed at low temperature. However, as we shall see, there is a temperature-dependent entropic gain associated to these lines due to these excitations, and this entropic gain eventually becomes larger than the entropic cost of a simple DDW. As a consequence, the free energy of a C line decreases as the temperature is raised up to a critical temperature where it becomes negative. Above this critical temperature a finite density of C lines develops, driving a phase transition where the antiferromagnetic long-range order of the dense rows is destroyed.
@@
Topological devil’s staircase. Numerical investigation using the Corner Transfer Matrix Renormalization Group (CTMRG) algorithm [41, 44, 45, 46] (see EM2) confirms that a transition where defect lines condense indeed takes place, proving that the entropy gain is sufficient to overcome the energy cost of defects (and the entropy loss of the zero energy DDW in the first place). In order to detect the transition, we use that $n_{A}$ (density of A lines) is equal to the density of crossed hexagons (in the dual bond representation), while $n_{C}$ (density of C lines) is the density of ferromagnetic nearest-neighbor pairs along a dense row. Thus, both $n_{A}$ and $n_{C}$ are written as local observables which can be easily calculated using CTMRG.
@@
Let us look in more details at what happens above the transition. If it was a usual Kasteleyn transition, the density of C-lines should grow continuously. However, this is clearly not the case, as shown in Fig. 2: this density jumps to a strictly positive value at the transition $T/J_{2}=1.78$. At the same time, the density of A-lines jumps to a value strictly smaller than 1/3, and most remarkably the ratio of the density of C-lines to that of A-lines is rigorously equal to 1. The interpretation is quite clear: since the two types of lines must avoid each other, C-lines must appear between pairs of A-lines. The configuration realized just above the transition corresponds to exactly one C-line between two consecutive A-lines, while the density adapts to a value different from 1/3 to minimize the total free energy. Looking at system snapshots sampled using CTMRG [41], shown in Figs. 3 (a) to (c), the constancy of the number of C lines between pairs of consecutive A lines was seen to hold at all temperatures. It is a consequence of the effective repulsion between adjacent C strings, due to the fact that excitations along consecutive C lines (not separated by an A) constrain each other.
@@
Figure 3: Topological plateaus are shown for increasing temperatures.(a-c): snapshots sampled from the CTM environment ($\chi=80)$ for (a) $p=n_{C}/n_{A}=1$ ($T=3J_{2}$) (b) $p=2$, $T=3.4J_{2}$ (c) $p=3$, $T=3.8J_{2}$ plateaus, with C-lines and green decorations. The number $p$ of C-lines between A strings is easily evaluated as the number of vertical orange bonds crossing a dense row of spins. Second row (d-f): corresponding magnetic structure factors exhibiting $n_{C}/n_{A}+2$ peaks at distances $2\pi/\bar{l}_{1}$. The structure factors are obtained by averaging over more than 2700 snapshots of size $40\times 40$.
@@
Now, we know from the case of the triangular lattice that the restoration of the $\mathbb{Z}_{3}$ symmetry requires the formation of a 2D graph where C-lines can develop simultaneously in all directions, to form domains with different orientations of arrows. This can only occur if the A-lines have completely disappeared. The transition into the high temperature phase must therefore correspond to the vanishing of the density of A-lines. Since the density of C-lines must remain finite, this implies that the ratio of the density of C-lines to that of A-lines diverges at the transition. With the (numerically validated, Fig. 3 a to c) assumption that the number of C-lines is the same between all pairs of consecutive A-lines, this implies that this ratio takes integer values from 0 ($\mathbb{Z}_{2}$ ordered phase) to infinity (high temperature phase). This is supported by our numerical data, where the phases up to a ratio equal to 4 could be unambiguously resolved. Higher plateaus are beyond the reach of our numerics [41].
@@
We are left with the remarkable conclusion that there is an infinite sequence of first-order phase transitions before the $\mathbb{Z}_{3}$ symmetry is restored, at which all physical quantities (energy, densities of domain walls, order parameters) jump. This is reminiscent of the devil’s staircase observed in the ANNNI and other models where the wave vector, which in principle could change continuously with temperature, actually jumps between commensurate values until the high-temperature phase is reached. The mechanism at play in our case is, however, qualitatively different. Indeed, the density of strings, hence the average distance between domain walls or the wave-vectors of the intermediate phases, are not constant inside the plateaus (see insets of Fig.2c). They change continuously in a narrow interval whose width is however much larger than our numerical resolution, so there is no doubt that the wave-vector is not pinned to a commensurate value. This devil’s staircase is thus not related to commensurability. It is actually of topological origin: it is the number of C domain walls between consecutive A domain that changes in steps of 1 upon increasing the temperature.
@@
Discussion. To conclude, let us briefly discuss the implications of the present results. The basic ingredient of the topological devil’s staircase reported here is the existence of two types of line defects in a constrained model, one that has a finite density in the ground state, the other one that can only develop at a finite temperature by an entropic mechanism, as in the Kasteleyn transition. With the additional assumption that these lines repel each other, a standard one for entropically induced defect lines, the conclusion that the transition takes the form of a devil’s staircase, in which the number of additional lines between consecutive ground state ones takes all integer values, is quite generic. Indeed, one can show that a statistical physics model of line defects based on these simple assumptions is able to reproduce the main features of the constrained Ising model studied here. One can also show that the same physics is present in a simple extension of 1D Hubbard model with two types of fermions, the role of the temperature being played as usual by the chemical potential. Finally, the present results also have very interesting consequences for the more realistic model with finite $J_{1}$ and $J_{3}$. As long as $J_{2}$ is not too large, the first transitions turn into very sharp crossovers, with in particular a remarkable series of sharp peaks in the low temperature specific heat. All these extensions will soon be reported elsewhere. In a broader context, considering recent progress in exploring and understanding constrained models [50, 51], we may expect to find other exotic transitions related to the Kasteleyn mechanism [52]. This exotic “classical” physics may also form a fertile ground for quantum models [51, 53, 54, 55].
@@
EM1 Mapping to strings
We show how to map the configurations of the constrained kagome Ising antiferromagnet with infinite $J_{1}$ and $J_{3}$ interactions onto the trajectories of three kinds of domain walls. We first note that the minimization of $J_{1}$ and $J_{3}$ interactions imposes local constraints on the kagome star clusters, allowing only the five classes of local configurations listed in Fig. EM1. As stated in the main text, each triangle is crossed by exactly one dual lattice bond, meeting at least one other bond at the center of a hexagon to form an arrow. The key is that arrows form domains with identical orientations, enabling the identification of the strings as domain walls. In practice, we map the kagome Ising configurations onto six-state clock model configurations by attributing to each triangle of the kagome lattice the direction of the arrow piercing it (see Fig. EM1 for the required conventions). This mapping is not surjective: the ground-state rules of the strings phase impose constraints on the allowed clock model configurations.
@@
  • III: Walls between domains with directions that differ by 60 or 120 degrees, in purple.
Types I and II directly correspond to the A and B lines used to describe the partially ordered family of states of the kagome strings phase, while C-lines are combinations of type III DWs.
@@
The partition function of a classical lattice model can be calculated as the contraction of a tensor network (TN), an idea directly related to transfer matrix studies of 2D Ising magnets [56, 57, 58]. The exact contraction of a two-dimensional TN is exponentially hard. We use an approximate contraction algorithm, the Corner Transfer Matrix Renormalization Group (CTMRG) [59, 60, 44]. In CTMRG, the infinite environment surrounding a site is approximated by the product of corner ($C_{i}$) and edge ($E_{i}$) transfer matrices, truncated to a finite cutoff bond dimension $\chi$, as shown in Fig. EM2a. In order to calculate the optimal approximation for the environment in the thermodynamic limit, tensors are iteratively added to the network and renormalized into the corner and edge tensors until a fixed point is reached. Truncation to $\chi$ limits long-range correlations, inducing a finite correlation length even in critical systems. Once the iteration procedure converges, the expectation value of local observables are obtained from exact TN contractions.
@@
Figure EM2: Tensor network. (a) CTRMG idea: the environment surrounding a cluster is approximated by corner ($C_{i}$) and edge ($E_{i}$) tensors, which are truncated to a cutoff dimension $\chi$. We use single-site directional CTMRG [61] with the isometries introduced in Ref. [62]. (b) Each tensor in the network contains the Bolzmann weight of a kagome star [45]. Contraction of the legs of the tensors impose the matching of spin configurations on overlapping sites. (c) If configurations related by a global spin-flip symmetry are identified, there are 4 different states for each triangle. We represent them by drawing a dual lattice bond between adjacent spins pointing in the same direction. The last one is not allowed in the constrained limit ($J_{1},J_{3}\rightarrow\infty$).

### v2

@@ Abstract
Two famous examples of unusual phase transitions are particularly relevant here. The first is the Kasteleyn transition, originally described by Kasteleyn in 1963 [7] in the context of dimer models [8, 9]. Also known as the Pokrovsky-Talapov transition, it occurs in constrained models [10, 11, 12] where the only defects that respect the constraints span the whole lattice; thus, their energy grows linearly with the system size. As a consequence, they are completely absent up to the temperature at which their free energy becomes negative due to the entropic contribution.
In frustrated models, the Kasteleyn transition may appear when constraints are imposed by infinite coupling constants. It is well known to play a key role in kagome ice [13, 14, 15], as well as in spin ice under a [100] field [16, 17, 18, 19, 20]. A case of direct relevance for the present study is the antiferromagnetic Ising model on the triangular lattice with second neighbor coupling [21]. The ground state consists of alternating lines of up- and down spins. It breaks the $\mathbb{Z}_{3}$ rotational lattice symmetry and the $\mathbb{Z}_{2}$ spin-flip symmetry. If a fifth-neighbor coupling is included, and in the constrained limit of infinite first-neighbor coupling, the $\mathbb{Z}_{2}$ symmetry is first restored through a Kasteleyn transition before the $\mathbb{Z}_{3}$ symmetry is restored through a first-order transition.
@@
In this Letter, we study the antiferromagnetic Ising model on the kagome lattice with up to third neighbor couplings (Eq. 1) in a phase that breaks the rotational $\mathbb{Z}_{3}$ symmetry and the $\mathbb{Z}_{2}$ spin-flip symmetry. In the appropriate constrained limit of infinite first and third neighbor couplings, where the only allowed defects are system-spanning, we show that the transition is a highly exotic form of a Kasteleyn mechanism giving rise to a topological devil’s staircase[25, 26]. Although linear defects are completely suppressed at low temperature, as in the Kasteleyn transition, their density, instead of growing as $(T-T_{c})^{1/2}$, grows through an infinite and increasingly dense sequence of jumps. By contrast to the standard devil’s staircase, this behavior is not related to commensurate steps, but is controlled by an integer index of topological origin, the ratio of the densities of linear defects and of zero-energy domain walls.
Figure 1: Spin configurations of the kagome lattice Ising antiferromagnet Eq. (1) are represented by blue and red dots. Orange segments show an alternative representation. (a) Main family of strings ground states : domains of left- and right-pointing orange arrows are delimited by A and B strings. The first dense and sparse rows which make up the Kagome lattice are highlighted as solid or dashed gray lines. Inset: labeling of the isotropic Ising couplings. (b) Zero-energy double domain walls (DDWs) correspond to arrows rotated all in the same direction and are entropically suppressed in the thermodynamic limit. A thick purple line highlights the center of each DDW. (c) C-string excitations in the first plateau are characterized by different orientations of the arrows inside the DDWs, which introduce an energy cost. The density of C-strings ($n_{C}$) is equal to the density of ferromagnetic nearest-neighbors on a horizontal dense row.
Model and partially ordered ground states. We focus on the properties of the constrained model that corresponds to $J_{1}$ and $J_{3}$ infinite in the Hamiltonian:
@@
where $\langle i,j\rangle_{n}$ denote pairs of $n^{\mathrm{th}}$ nearest-neighbors on the kagome lattice, and $\sigma_{i}=\pm 1$. In contrast to already extensively studied models where the third-neighbor interaction is only across the hexagon [27, 28, 29, 30, 31, 32] or along the nearest-neighbor bonds [33, 34, 35, 36, 37, 38, 39]  [^1] A notable case is the SM of Ref. [71], which studies the $J_{2}=0$ limit with both 3rd neighbor couplings being non-zero., here, the third-neighbor interaction is set to the same value in both directions, forming triangular sublattices (Fig. 1a). For finite $J_{1}$ and $J_{3}$, the range $0<J_{2}\leq J_{3}<J_{1}$ corresponds to a single phase of ground state energy per site $E/N=-2J_{1}/3+2J_{2}/3-J_{3}/3$, referred to as the strings phase in Ref. [41]. In that phase, the macroscopic degeneracy of the nearest-neighbor model is only partially lifted, yielding a residual entropy equal to 1/3 of that of triangular lattice antiferromagnet.
@@
This entropy is compatible with a simple family of ground states, illustrated in Fig. 1(a). Seeing the kagome lattice as a stacking of dense and sparse rows, the dense rows form a rectangular lattice. Consider a configuration on this rectangular lattice where the order is antiferromagnetic in the direction of the dense rows and ferromagnetic in the perpendicular direction. The energy minimum is reached if the remaining spins, which live on a $J_{3}$ triangular sub-lattice, are in the ground state of that model, with one frustrated bond on each $J_{3}$ triangle. Since the number of these spins is 1/3 of the total number of spins, the residual entropy of this family is 1/3 of that of the triangular lattice. To visualize these states, it is convenient to draw dual-lattice bonds through the unsatisfied $J_{1}$ bonds. In the ground state, these dual bonds must not appear isolated; instead, they must belong to a pair of bonds meeting at 60º in the center of a hexagon, forming an arrow. If we define domains as contiguous regions where all arrows point in the same direction, the partially ordered family can be constructed by partitioning the lattice into domains with antiparallel directions  [^2] This is similar in spirit to the construction of the triangular lattice Ising antiferromagnetic ground states starting from the stripe order and introducing double domain walls, see e.g. [21, 72] and references therein.. The boundaries between the uniform regions form domain walls living on the bonds of the triangular lattice and connecting the centers of the hexagons. We call these domain walls A or B lines if they are surrounded by converging or diverging arrows, respectively (see EM1), or equivalently if they go through crossed or empty hexagons. In order to minimize $J_{3}$ interactions, the A and B lines may never propagate in parallel to the arrow directions or intersect, which forces them to form system-spanning zero-energy domain walls. Note that the number of A and B lines may at most differ by one since every domain must be bounded by an A line on one side and a B line on the other. In the thermodynamic limit, their densities must therefore be equal.
@@
In this partially ordered family of states, the trajectories of A and B lines determine the spin configuration on the disordered $J_{3}$ triangular sub-lattice, where they separate antiferromagnetically aligned spins in the same row. Since adjacent spins on sparse rows have a $2/3$ probability of being antiferromagnetically aligned, the average density of A and B strings in the thermodynamic limit is $1/3$.
@@
Figure 2: Numerical results from CTMRG for $\chi=100$ suggest a topological devil’s staircase. Only temperatures where the results are converged in bond dimension are shown (for more details on the bond-dimension dependence of CTMRG results, see [43]). (a) Energy. (b) Ratio of the strings density, revealing well-defined, integer-valued plateaus. (c) Density of A and C strings, related to the $\mathbb{Z}_{3}$ order parameter (see EM3). A careful inspection reveals a continuous temperature dependence in the apparent plateaus of $n_{C}$, as shown in the two insets.
@@
Interestingly, there is another family of zero energy extended defects, namely double domain walls (DDW) formed by replacing a B line with any number of arrows tilted by 60 or 120 degrees, as shown in Fig. 1(b). Such a DDW creates a defect in the antiferromagnetic order of the dense rows, and if a finite density of them (with random positions) was present in the ground state the $\mathbb{Z}_{2}$ symmetry would not be broken. These DDWs may meander in the same way that A and B lines do, but they have a thickness of one lattice unit. Thus, the entropy of configurations with one DDW is equivalent to the entropy of configurations with only A and B lines in a smaller system where one column has been removed. Since entropy is extensive, removing a column has an entropic cost linear in system size, and therefore DDWs are completely suppressed in the thermodynamic limit, inducing an entropy-driven partial order [44, 45]. This is in agreement with the numerical observation [41] that the residual entropy is exactly that of the simple family of states with only A and B lines described above. In addition, there are also ground states which cannot be described by replacing B lines with DDW’s. Their number grows sub-extensively with system size, and therefore they also do not contribute to macroscopic properties of the ground state [43].
@@
Excitations in the constrained limit. Let us discuss more specifically the constrained limit of infinite $J_{1}$ and $J_{3}$. In that limit, the energy of the triangles formed by $J_{1}$ and $J_{3}$ interactions is minimized, leading to the local configurations listed in FIG. EM1. In this constrained model, the only possible excitations at finite temperature correspond to deviations from the $J_{2}$ ground state rule, which imposes a single bond to be unsatisfied in each $J_{2}$ triangle. Such excitations are only possible if a DDW is present (DDW which, in the constrained limit, must be system spanning; for a proof, see [43]). They are rooted in the development of domains with changes of the orientation of the arrows inside the DDW, possibly accompanied by portions of B-lines which decorate the DDW (see panel (c) of Fig. 1). From now on, these DDW with possible changes of orientation and decorations will be called C-lines. The energy cost of such C-lines comes from the presence of two types of local defects: (i) hexagons with two arrows that form a “K” instead of a cross, with energy $4J_{2}/3$, and (ii) hexagons with three arrows forming a “star”, with energy $4J_{2}$. As in the Kasteleyn transition, excitations with a finite density of defects are completely suppressed at low temperature. However, as we shall see, there is a temperature-dependent entropic gain associated to these lines due to these excitations, and this entropic gain eventually becomes larger than the entropic cost of a simple DDW. As a consequence, the free energy of a C line decreases as the temperature is raised up to a critical temperature where it becomes negative. Above this critical temperature a finite density of C lines develops, driving a phase transition where the antiferromagnetic long-range order of the dense rows is destroyed.
@@
Topological devil’s staircase. A numerical investigation using a directional version of the Corner Transfer Matrix Renormalization Group (CTMRG) algorithm [43, 46, 47, 48] (which can capture highly anisotropic correlations, see EM2) confirms that a transition where defect lines condense indeed takes place. Note that the density $n_{A}$ of A lines is equal to that of crossed hexagons (in the dual bond representation), while the density $n_{C}$ of C lines is equal to that of ferromagnetic nearest-neighbor pairs along a dense row. Both can thus be written as local observables, hence easily calculated using CTMRG.
@@
Let us look in more details at what happens above the transition. If it was a usual Kasteleyn transition, $n_{C}$ should grow continuously. However, this is clearly not the case, as shown in Fig. 2: $n_{C}$ jumps to a strictly positive value at the transition $T/J_{2}=1.78$. At the same time, $n_{A}$ drops to a value strictly smaller than 1/3, and most remarkably the ratio $n_{C}/n_{A}$ is rigorously equal to 1. The interpretation is quite clear: since the two types of lines must avoid each other, C-lines must appear between pairs of A-lines. The configuration realized in the first plateau just above the transition corresponds to exactly one C-line between two consecutive A-lines, as confirmed by snapshots sampled using CTMRG [43], a consequence of the effective repulsion between adjacent C strings since excitations along consecutive C lines (not separated by an A) constrain each other. Note that the densities themselves are slightly temperature dependent.
@@
Figure 3: Topological plateaus are shown for increasing temperatures.(a-c): snapshots sampled from the CTM environment ($\chi=80)$ for (a) $p=n_{C}/n_{A}=1$ ($T=3J_{2}$) (b) $p=2$, $T=3.4J_{2}$ (c) $p=3$, $T=3.8J_{2}$ plateaus, with C-lines and green decorations. The number $p$ of C-lines between A strings is easily evaluated as the number of vertical orange bonds crossing a dense row of spins. Second row (d-f): corresponding magnetic structure factors exhibiting $n_{C}/n_{A}+2$ peaks at distances $2\pi/\bar{l}_{1}$. The structure factors are obtained by averaging over more than 2700 snapshots of size $40\times 40$ and clearly highlight the anisotropy of correlations, also evident in the real-space plots [43].
@@
Now, the restoration of the $\mathbb{Z}_{3}$ symmetry requires the formation of a 2D graph where C-lines can develop simultaneously in all directions, to form domains with different orientations of arrows [21]. This can only occur if the A-lines have completely disappeared, hence if the ratio $n_{C}/n_{A}$ diverges at the transition. With the assumption, numerically validated (see Fig. 3 a to c), that the number of C-lines is the same between all pairs of consecutive A-lines, this implies that this ratio takes integer values from 0 ($\mathbb{Z}_{2}$ ordered phase) to infinity (high temperature phase). This is supported by our numerical data, where the phases up to a ratio equal to 4 could be unambiguously resolved. Higher plateaus are beyond the reach of our numerics [43].
@@
We are left with the remarkable conclusion that there is an infinite sequence of first-order phase transitions before the $\mathbb{Z}_{3}$ symmetry is restored, at which all physical quantities (energy, densities of domain walls, order parameters) jump. This is reminiscent of the devil’s staircase observed in the ANNNI and other models where the wave vector jumps between commensurate values until the high-temperature phase is reached. The mechanism at play in our case is, however, qualitatively different since the density of strings, hence the average distance between domain walls or the wave-vectors of the intermediate phases, are not constant inside the plateaus (see insets of Fig.2c). This devil’s staircase is thus not related to commensurability, but of topological origin: it is the number of C domain walls between consecutive A domain that changes in steps of 1 upon increasing the temperature.
@@
Discussion. To conclude, let us briefly discuss the implications of the present results. The core ingredient of the topological devil’s staircase reported here is the existence of two types of line defects in a constrained model, one that has a finite density in the ground state, the other one that can only develop at a finite temperature by an entropic mechanism, as in the Kasteleyn transition. With the additional assumption that these lines repel each other, a standard one for entropically induced defect lines, the conclusion that the transition takes the form of a devil’s staircase, in which the number of additional lines between consecutive ground state ones takes all integer values, is quite generic. Indeed, one can show that a statistical physics model of line defects based on these simple assumptions is able to reproduce the main features of the constrained Ising model studied here. Similarly to spin ice [18, 16], a simple quantum-one-dimensional analogue of the present mechanism would be to interpret one of those lines (C-strings) as bosonic worldines. In this analogy, the other type of particles (A) induces a finite-size confinement for the first ones, leading to the observed staircase of jumps as they are slowly driven apart by the increasing chemical potential. The same physics is also present in a simple extension of the 1D Hubbard model with two types of fermions, the role of the temperature being played as usual by the chemical potential. All these related models are reminiscent of the wide variety of origins of the Kasteleyn mechanism, related to many experimental realizations [12], and may open paths to observe this unusual series of transitions in magnetic oxides, artificial spin systems, or Rydberg atom arrays. Finally, the present results also have very interesting consequences for the more realistic model with finite $J_{1}$ and $J_{3}$. As long as $J_{2}$ is not too large, the first transitions turn into very sharp crossovers, with in particular a remarkable series of sharp peaks in the low temperature specific heat. All these extensions will soon be reported elsewhere. In a broader context, considering recent progress in exploring and understanding constrained models [52, 53], we may expect to find other exotic transitions related to the Kasteleyn mechanism [54]. This exotic “classical” physics may also form a fertile ground for quantum models [53, 55, 56, 57].
Data availability. The data supporting the findings of this article will be made openly available on 21 Jan. 2027 [58].
@@
EM1 Local constraints and mapping to strings
We show how to map the configurations of the constrained kagome Ising antiferromagnet with infinite $J_{1}$ and $J_{3}$ interactions onto the trajectories of three kinds of domain walls. The constraints imposed by the $J_{1},J_{3}\rightarrow\infty$ limit forbid ferromagnetic $J_{1}$ or $J_{3}$ triangles. As a consequence, spin configurations in each of the three triangular sublattices of $J_{3}$ bonds must be ground-states of the nearest-neighbor, triangular lattice Ising antiferromagnet. Intuitively, this leads to the extended defects description of the configurations in the constrained manifold: constraint-allowed configurations in each $J_{3}$ sublattice can be divided into topological sectors that can only be connected by non-local updates [21, 59, 60]. A more detailed explanation of the nature of these extended defects, and a proof that they must be system-spanning, is given in the SI [43]. Here, we focus on the definition of mapping from the spin configuration to the strings. The minimization of $J_{1}$ and $J_{3}$ interactions imposes local constraints on the kagome star clusters, allowing the five classes of local configurations listed in Fig. EM1. As stated in the main text, each nearest-neighbour triangle is crossed by exactly one dual lattice bond, meeting at least one other bond at the center of a hexagon to form an arrow. Arrows form domains with identical orientations, enabling the identification of strings as domain walls. In practice, we map the kagome Ising configurations onto six-state clock model configurations by attributing to each triangle of the kagome lattice the direction of the arrow piercing it (see Fig. EM1 for the required conventions). This mapping is not surjective: the ground-state rules of the strings phase impose constraints on the allowed clock model configurations.
@@
  • III: Walls between domains with directions that differ by 60 or 120 degrees, in blue.
The three types of domain walls are drawn in illustrative configurations in the bottom of Fig. EM1. Domain walls of types I and II correspond directly to the A and B strings whose trajectories describe the partially ordered family of states mentioned in the main text. Type III domain walls can be combined to form double domain walls (DDW’s), which surround columns of oblique arrows. We call each such column a C-string, drawn in purple.
@@
The partition function of a classical lattice model can be calculated as the contraction of a tensor network (TN), an idea directly related to transfer matrix studies of 2D Ising magnets [61, 62, 63]. The exact contraction of a two-dimensional TN is exponentially hard. We use an approximate contraction algorithm, the Corner Transfer Matrix Renormalization Group (CTMRG) [64, 65, 46]. In CTMRG, the infinite environment surrounding a site is approximated by the product of corner ($C_{i}$) and edge ($E_{i}$) transfer matrices, truncated to a finite cutoff bond dimension $\chi$, as shown in Fig. EM2a. The directional CTMRG algorithm [66] we use does not rely on any assumption of rotational or parity symmetries and has no issue capturing highly anisotropic correlations [67]. In the course of CTMRG, tensors are iteratively added to the network and renormalized into the corner and edge tensors until a fixed point is reached. Truncation to $\chi$ limits long-range correlations, inducing a finite correlation length even in critical systems.
@@
Figure EM2: Tensor network. (a) CTRMG idea: the environment surrounding a cluster is approximated by corner ($C_{i}$) and edge ($E_{i}$) tensors, which are truncated to a cutoff dimension $\chi$. We use single-site directional CTMRG [66] with the isometries introduced in Ref. [68]. (b) Each tensor in the network contains the Boltzmann weight of a kagome star [47]. Contraction of the legs of the tensors impose the matching of spin configurations on overlapping sites. (c) If configurations related by a global spin-flip symmetry are identified, there are 4 different states for each triangle. We represent them by drawing a dual lattice bond between adjacent spins pointing in the same direction. The last one is not allowed in the constrained limit ($J_{1},J_{3}\rightarrow\infty$).